two executive in cities 300 mi apart drive to a business meeting at a location between their cities. they met after 4 hours. find the speed of each car if one travels 20 miles per hour faster than the other
Answer by amoresroy(361)

Let x = speed of slower car
x+20 = speed of faster car
4(x +x+20) = 300
8x + 80 = 300
8x = 220
x = 27.5
x+20 = 27.5+20 = 47.5
The speed of each car is 27.5mph and 47.5 mph.